OpenStudy (anonymous):

is it possible to find the zeros of this function: f(x) = 2x-2e^(-2x) Without the aid of a graphing calculator to do it for you?

OpenStudy (anonymous):

hmm f(1) = 2(1)-2e^(-2(1)) f(1) = 2-(2e^(-2)) I don't think that would equal to zero...

OpenStudy (anonymous):

i really don't mind, I just want to know if it is possible to do this without a graphing calculator (without using the "intersect" function).

OpenStudy (anonymous):

sorry I meant "zero" function

OpenStudy (anonymous):

you mean f =0?

OpenStudy (anonymous):

yeah, like y = 0

OpenStudy (reemii):

It's not possible to get it by a formula.

OpenStudy (anonymous):

ok so algebraically its impossible?

OpenStudy (reemii):

I know you can't find the solution of \(e^{-x}=x\) so I bet it's the same for \(e^{-2x}=x\), which is what you have since you want to find the zero of \(2(x-e^{-2x})\).

OpenStudy (anonymous):

Ah ok then. I was just curious to see if this was possible but I guess not. Thanks anyways!